If one side of a square is increased by 2 inches and an adjacent side is decreased by 2 inches, the area of the resulting rectangle is 32 square inches. Find one side of the square.
Answer by checkley77(12844)

(x+2)(x-2)=32
x^2+2x-2x-4=32
x^2-4-32=0
x^2-36=0
(x+6)(x-6)=0
x-6=0
x=6 the length of the original side of the square.
Proof:
(6+2)(6-2)=32
8*4=32
32=32